OpenStudy (kevin):

If z is the smallest positive integer giving remainder 5 when divided by 13 and giving remainder3 when divided by 18, what's the remainder when divided by 11?

OpenStudy (kevin):

@zepdrix @Nnesha @.Sam. @jiteshmeghwal9 @Frozen15 @legomyego180 @karim728

OpenStudy (jiteshmeghwal9):

2 ??

OpenStudy (jiteshmeghwal9):

Is 2 the answer ?

OpenStudy (kevin):

yes

OpenStudy (kevin):

can u tell me how is it?

OpenStudy (jiteshmeghwal9):

I solved it thru hit & trial method

OpenStudy (kevin):

there is no formula?

OpenStudy (kevin):

from here : z = 13a + 5 = 18b + 3 we should use hit & trial method?

OpenStudy (jiteshmeghwal9):

z=13x+5 z=18p+3 => 13x=(z-5) 18p=(z-3)

OpenStudy (jiteshmeghwal9):

from here i used hit & trial method

OpenStudy (kevin):

can u explain me how is it?

OpenStudy (jiteshmeghwal9):

13*3=39 18*2=36 these 2 multiples do not satisfies the equations 13*4=52 18*3=54 these 2 multiples satisfies the equations i made. The rule i used in choosing these certain multiples is tht there should not be any large difference between the multiples.

OpenStudy (jiteshmeghwal9):

let's see if any user knows a formula to solve the question. By the way question is very nice

OpenStudy (kevin):

thx

OpenStudy (kevin):

I still have more question. But it's very long question. If u don't mind anyway? @jiteshmeghwal9

OpenStudy (jiteshmeghwal9):

plz

OpenStudy (jiteshmeghwal9):

where is the question ?

OpenStudy (kevin):

check my question